OpenStudy (anonymous):

Giving Medals!!!! Solve for x: |x − 2| + 10 = 12

OpenStudy (anonymous):

x=-1

OpenStudy (anonymous):

|x-2|=12-10=2 |x-2|=2 on removing mode \[x-2=\pm2\]

OpenStudy (anonymous):

x-2=2 and x-2=-2 x=2+2 and x=-2+2 x=4 or x=0

OpenStudy (anonymous):

once you have your x which is -2 you do letters to the left and numbers to the right. once you do that you should have -2x=12-10 since 10 is before the = you move it behind and it becomes -10. Once you do that you should get -2x=12-10 which = -2x=2 then you divide by -2 and you should get -1. So it's x=-1